OpenStudy (anonymous):

Can anyone show me how to multiply this fraction with all the steps? 3/4 x1/8 =

OpenStudy (swissgirl):

\( \huge \frac{3} {4} * \frac{1} {8}\) So we mutiply the 2 numerators and the 2 denomenators: \(\huge \frac{3*1} {4*8}=\frac{3} {32}\)

OpenStudy (anonymous):

Hi, how are you doing today? Thanks for responding. Is that all I have to do with this problem? That is what I did and that's the answer, I came up with... but I wasn't sure I was right.

OpenStudy (swissgirl):

Yes that is all you have to do. Well you may have to reduce to lower terms which we cant in this case. So the answer is what I have shown u above

OpenStudy (anonymous):

Great!! Thanks. i will close this out now. I understand this one. Thanks for helping me.

OpenStudy (swissgirl):

non problema